OpenStudy (anonymous):

Find the product of (3x − 7y)2. 9x2 − 42xy + 49y2 9x2 + 42xy + 49y2 9x2 − 49y2 9x2 + 49y2

Parth (parthkohli):

\(\Large \color{Midnightblue}{\rightarrow (a - b)^2 = a^2 - 2ab + b^2 }\)

Parth (parthkohli):

This above thing is called an identity.

OpenStudy (anonymous):

A thing to go by, right?

Parth (parthkohli):

Yes, this is basically a general formula that applies to all numbers.

OpenStudy (anonymous):

So, it would be, 9x^2 - 42 + 49.

Parth (parthkohli):

9x^2 - 42xy + 49. If this was a mistake, then yes.

OpenStudy (anonymous):

Why would it be 9x^2 - 42xy + 49y^2?

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

okay, so.. Find the product of (4x + 3y)(4x − 3y).

Parth (parthkohli):

^^ \((a + b)(a - b) = a^2 - b^2 \)

OpenStudy (anonymous):

Just give me the formula and I can go by it.

OpenStudy (anonymous):

16x^2 - 9y^2??

Parth (parthkohli):

Okay, what all formulae?

Parth (parthkohli):

\(\Large \color{Midnightblue}{\rightarrow a^2 - b^2 = (a + b)(a - b) }\)

OpenStudy (anonymous):

Formulas are things to go by, like an identify. I call them formulas. The only problem I see and know that I have, is that; I don't know which formula to go by for each problem? you know what I'm saying?

OpenStudy (anonymous):

16x^2 - 9y^2??

Parth (parthkohli):

I know what formula is!

OpenStudy (anonymous):

Why are you yelling at me?

OpenStudy (anonymous):

I know I'm dumb, but dang.

Parth (parthkohli):

\(\Large \color{Midnightblue}{\rightarrow 16x^2 - 9y^2 = (4x + 3y)(4x - 3y) }\)

Parth (parthkohli):

@careless850 All learners are lol, but when you get the concepts, then it is a piece of cake.

OpenStudy (anonymous):

I know. Like this one Find the product of (x + 3)2. we just did one just like it, but I already forgot what formula to go by..

OpenStudy (anonymous):

Find the product of (x + 3)^2.

Parth (parthkohli):

Okay, \(\Large \color{Midnightblue}{\rightarrow 1) (a \pm b)^2 = a^2 \pm 2ab + b^2 }\)

OpenStudy (anonymous):

Should I write them down?

Parth (parthkohli):

\(\Large \color{Midnightblue}{\rightarrow 2) a^2 - b^2 = (a + b)(a - b) }\)

Parth (parthkohli):

If you want to.

OpenStudy (anonymous):

ALright..

Parth (parthkohli):

\(\Large \color{Midnightblue}{\rightarrow (a + b)^3 = a^3 + b^3 + 3ab^2 + 3a^2b}\)

OpenStudy (anonymous):

So, on that problem the answer is, x^2 + 6x + 9?

Parth (parthkohli):

It is.

OpenStudy (anonymous):

Find the product of (3x + 7y)2.

OpenStudy (anonymous):

I would use... 1.

Parth (parthkohli):

Sure!

Parth (parthkohli):

Right.

OpenStudy (anonymous):

So, 9x^2 + 42 + 49y^2?

OpenStudy (anonymous):

Why do I have to put 42xy?

OpenStudy (anonymous):

THAT I don't understand..

Parth (parthkohli):

What is \(2(3x)(7y)\)?

OpenStudy (anonymous):

The answer is 9x^2 + 42xy + 49y^2.

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

42xy.

Parth (parthkohli):

Then? :P

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

The product of (a + b)(a − b) is a2 − b2.

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

The product of (a + b)(a − b) is a^2 − b^2. is The product of (a + b)(a − b) is a2 − b2. Sometimes Always Never

OpenStudy (anonymous):

Omg.

Parth (parthkohli):

Always!

OpenStudy (anonymous):

The product of (a + b)(a − b) is a^2 − b^2. Sometimes Always Never

Parth (parthkohli):

Identity works for each and every number.

OpenStudy (anonymous):

Hmm.. Okay. So, Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

(x + 4)(x - 4) = x^2 - 16 (x + 4)(x + 4) = (x + 4)^2

OpenStudy (anonymous):

Okay. Compare and contrast, in complete sentences, the similarities and differences of the two.

Parth (parthkohli):

I don't know, can you do that please?

OpenStudy (anonymous):

Yes.